# Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination

### v2

@@ 1 Introduction
Visual realism with global illumination is a key demand of immersive virtual reality (VR) applications to intrigue users’ immersion arousal and respond realistically to virtual environments and events [47]. However, realistic rendering of such globally-illuminated scenarios in real time remains a big challenge, especially in handling complex lighting situations like dynamic changes of illumination and environment lighting [31]. While photo-realistic rendering techniques (e.g., Monte Carlo path tracing) can synthesize images with global illumination, they are often extremely time-consuming and may take hours to render. Meanwhile, various works attempt to use approximations along with classical rasterization-based rendering, thus approximating global illumination effects at interactive or real-time performance. They generally operate in either image or 3D space and adopt different precomputation techniques to reduce runtime cost [38, 22]. Recently, learning-based techniques for photo-realistic rendering were proposed, some of which combine data-driven methods with classic rendering pipeline, resulting in improved image quality or rendering speedup [6, 44]. These techniques, namely neural rendering [42], usually provide stronger supervision to generative models by applying the prior knowledge in classical rendering techniques to neural networks.
Predicting indirect illumination is, nonetheless, a challenging task. Prior works generally regard this as a special case of image-to-image translation, and use convolutional neural networks (CNN) for indirect/global illumination prediction. The auxiliary geometry features are provided as inputs to predict global illumination, further conditioned on input the direct lighting [32, 46] or emission maps [6]. However, several major challenges remain:
@@
To address these issues, we adopt a learning-based approach to address the issue of global illumination, following the principles of the rendering equation [20]. Our approach includes a learnable module that aggregates global features while utilizing geometry supervision. This module incorporates a modified multi-head attention mechanism [43] to calculate the distribution based on geometry features, which guides the global aggregation of spatial features. This non-local process can be analogized to an integral over screen space, where the geometry-aware attention works as a learnable counterpart of the geometry term. Our proposed module overcomes the network bottleneck of the deep stack of convolutional layers, requiring less computation and fewer parameters while better capturing the long-range dependencies. Moreover, it can be applied to other tasks where geometry supervision is available and long-range dependency modeling is preferred.
Second, we present a monochromatic model design, where each RGB channel is independently processed using a shading generator conditioned on shared features from a geometry encoder. The redundant inter-channel interactions (introduced by the interconnections of convolutional filters) can be eliminated, and this relieves the burden of training data and network capacity, and enforces the network to learn the accurate color of indirect light. Furthermore, this makes our model immune to color variations, therefore allowing better generalization to novel and complex scenes and dynamic lighting, which are not present in the training data. Since HDR images can provide a much wider range of brightness and contrast for the high sensitivity of the human visual system to enhance the user’s perception of immersion and presence in the immersive virtual environment [7, 28], our model is trained to synthesize HDR global illumination.
Finally, we created a novel synthetic HDR dataset of indoor scenes for network training based on the public indoor layout dataset 3D-Front [11]. Our dataset comes with well-annotated lighting, texture, and geometry features; the surfaces are also assigned with physically based materials, providing sufficient diversity for geometry features and visual appearances. We design an adversarial training framework to directly optimize our network on HDR data, where the discriminator and the perceptual loss are jointly used for network optimization.
@@ 2.1 Real-Time Global Illumination
Numerous works seek to acquire global illumination effects at a low time budget. Classical methods adopt different approximations to reduce performance costs. Some of them operate in 2D screen space, such as screen space ambient occlusion (SSAO) [1] and screen space directional occlusion (SSDO) [35], using rasterized direct lighting and g-buffer features to approximate inter-reflections between surfaces. While some others operate in 3D space, usually achieving more realism with higher computational costs and storage, such as voxel cone tracing (VXGI) [5] and light propagation volumes (LPV) [21]. Lastly, precomputation-based methods like precomputed radiance transfer (PRT) [38] reduce the runtime cost by computing pre-integrations on the environment and BRDFs. Recently, with the advances in deep learning, data-driven methods have also been adopted to improve rendering quality efficiently. For example, training CNNs to denoise low sample-count images [3], or predicting the proper sample count for real-time raytracing [26].
@@ 2.2 Image-to-Image Translation
The task of image-to-image translation converts images from one domain to another, e.g., image style translation, colorization, etc. The architecture of convolutional neural networks (CNN) has been the most commonly used technique for image-to-image tasks, such as U-Net [36] and ResNet [15]. After the generative adversarial network (GAN) proposed by Goodfellow et al. [13], it has been further extended to conditional generative models and introduced to image-to-image tasks as well [18]. Recently, there has been another trend of using pre-trained models (e.g., VGG-19 [37]) to capture structurally correlated high-level image features for network optimization, namely the perceptual loss [19]. We show that a model’s output quality and robustness could be enhanced by jointly introducing adversarial loss and perceptual loss into network optimization.
@@ 2.3 Network Conditioning
Image-to-image translation could also be regarded as conditional synthesis, where networks learn to generate images conditioned on input geometry or semantic labels. Most of the early works simply provide conditioning data as the input of the network [18]. In contrast, many recent works focus on finding better ways to condition the image generation process, e.g., using a depth map to explicitly guide the convolution [45], using learned transformations to modulate convolutional features [33]. While these works have achieved great success, they are mainly used along with convolution layers, and thus remain local conditioning methods.
Inspired by the great success of attention mechanisms in modeling global dependencies [43, 48] and motivated by our task formulation, we propose a feature aggregation method that models long-range dependencies guided by geometry features, where the aggregation weights are obtained from a modified multi-head attention mechanism. Different from the prior work on integrating attention to CNN [48], we 1) explicitly use geometry features to guide the aggregation of global features; then 2) extend it into multi-head attention [43], and find it sufficient to equip only bottleneck layers with attention. Both bring a significant reduction of computational overhead, with no observable performance degradation in our experiments.
@@ 2.4 Global illumination prediction
Several recent works use neural networks to predict global illumination given auxiliary shading and geometry features. Nalbach et al. [32] first introduced CNN into global illumination prediction. They use a U-shaped CNN, taking pixel-wise features as inputs and predicting globally illuminated images. Bi et al. [2] train GANs to predict global illumination, and Dai et al. [6] use the perceptual loss for network optimization. Then Xin et al. [46] used a lightweight CNN and predicted the indirect component similar to the task setup in our work. However, Nalbach et al. [32] used global illumination computed in screen space as their ground truth, and Dai et al. excluded textures in their ground truth data generation. Most prior works used LDR data for training or took no special care in training in linear HDR space. Moreover, they generally use stacked convolutional operations for shading learning, limiting the receptive fields to local neighborhoods, and thus it is difficult to predict global effects like indirect illumination.
@@ 3.1 Task Formulation
We aim to predict indirect lighting given direct lighting and auxiliary geometry features, including surface normal, depth, reflectance, etc. All these inputs can be efficiently computed using a rasterization-based rendering pipeline in real time. For simplicity, we assume all Lambertian surfaces within the scene, i.e., the BRDF term $f(\mathrm{p},\omega_{i},\omega_{o})$ is a constant for arbitrary $\omega_{i}$ and $\omega_{o}$. We first briefly review the concepts of rendering equation [20] and physically based rendering; by splitting the incident radiance $L_{i}$ into direct and indirect terms, we derive an alternative form of the surface rendering equation:
@@
where ${L}_{\mathrm{o}}^{\mathrm{d}}$ and ${L}_{\mathrm{o}}^{\mathrm{ind}}$ are the direct and indirect components of the scattering radiance, respectively. $G(\mathrm{p},\,\mathrm{p^{\prime}})=\frac{V(\mathrm{p^{\prime}}\leftrightarrow\mathrm{p}){\cos\theta\cos\theta^{\prime}}}{\parallel\mathrm{p}-\mathrm{p^{\prime}}\parallel}$ is the geometry term where $V(\mathrm{p^{\prime}}\leftrightarrow\mathrm{p})$ indicates the visibility between two vertices. We can obtain the direct lighting term ${L}_{\mathrm{o}}^{\mathrm{d}}$ by integrating over all light sources or by real-time alternatives [38, 10, 16]. We then feed ${L}_{\mathrm{o}}^{\mathrm{d}}$ into the network and predict the indirect component ${L}_{\mathrm{o}}^{\mathrm{ind}}$.
To apply fine-grained supervision to the network, we first decompose the direct and indirect components of the final illumination, letting the network focus on predicting the indirect component instead of global illumination. We then use reflectance factorization to demodulate shading and diffuse reflectance, thus effectively tasking the network with learning the demodulated shading (instead of scattered radiance), which has a relatively low frequency after multiple bounces within the scene. The final rendering is obtained by multiplying the predicted shading with the reflectance. Moreover, we propose a feature aggregation module for better modeling global dependencies, which is intuitively similar to an integral over screen space. We employ the modified attention mechanism [43] to impose the effects of distance and visibility on network models, which intuitively works like a learnable counterpart of the geometry term, as is discussed later.
@@ 3.2 Model Design
Overview Our model contains a shading generation branch and a geometry encoding branch, both using a U-shaped network with skip-connections [36]. We integrate 1) our feature aggregation module into the shading generator to model long-range dependencies; 2) the monochromatic design for a more consistent deep rendering process. We also utilize conditional modulation techniques [33] for better geometry-aware supervision. The overview of our network architecture is shown in Fig. 2.
Geometry Encoder We use an additional geometry encoder that encodes geometry properties (i.e., positions $\mathbf{P}$, normals $\mathbf{N}$, and depth $\mathbf{D}$), leaving the other inputs to the shading encoder. The encoded geometry features are provided to condition the synthesis process of the shading generator. The separation of the encoding process splits the task into spatial structure modeling and light transport simulation, thus alleviating the requirements on training data and network complexity. Instead of simply concatenating the geometry features to the inputs of the generator layers, we use a more effective method, namely geometry-aware conditional modulation (GCM) similar to [33]. Specifically, the conditional geometry features are transformed into scaling and biasing factors through learnable convolution layers as:
@@
Geometry-Aware Feature Aggregation Fully convolutional networks (FCN) will result in a limited-sized receptive field [46, 32, 6], i.e., the output pixel value is conditioned only on a local region. While deep stacks of convolutional operations can enlarge the receptive field, they also bring prohibitive computational costs, making network optimization difficult. In this work, we expect the network to generate indirect illumination from features of all locations within the image space. A learnable module is needed that can efficiently aggregate global features guided by geometry.
We use a modified multi-head attention mechanism [43] to acquire the weights for the feature aggregation module based on geometry features (i.e., positions, depth, and surface normals). For each spatial location, the module first calculates the attention weights based on the encoded geometry features $f_{g}$, then obtains the response of that position as a weighted combination of the features from all locations. The attention weights $\mathrm{Attn}(j\,|\,i)$ for location $i$ are computed as:
@@
In addition, we find it beneficial to use multiple learnable linear transformations within attention computation, i.e., obtain multiple attention distributions $\mathrm{Attn}_{k}$ using different sets of linear transformations $Q_{k}$, $K_{k}$, and $V_{k}$, where k=1hk=1...h is the index of attention head and $h=8$ in our experiments. The aggregated features $\hat{x_{k}}$ are then concatenated to produce the final output $\hat{x}$, resulting in the multi-head version [43] of the aggregation module. This allows the network to simultaneously attend to different subspaces of the spatial features, resulting in a better capacity for modeling long-range dependencies.
@@ 3.4 Adversarial Training and Loss Function
where $\omega_{\mathrm{c}}$, $\omega_{\mathrm{p}}$, and $\omega_{\mathrm{a}}$ are the hyperparameters controlling the weight of each loss term, we use empirical values where $\omega_{\mathrm{c}}=0.7$, $\omega_{\mathrm{p}}=0.28$, $\omega_{\mathrm{a}}=0.02$ in our experiments. We use smooth L1 loss [34] as our content loss. While the L1 loss could not capture the inter-pixel correlations or the structural information of the image, we thus additionally introduced a PatchGAN [18] as the discriminator to obtain the adversarial loss term, which helps recover high-frequency local details. Moreover, it is an ill-posed problem to predict indirect illumination given direct lighting in screen space only (instead of in 3D solutions). While classic pixel-wise loss (e.g., L1, L2) generally tends to predict the average value over all possible solutions [27], we hope the adversarial training can guide the generator to approximate one particular solution, thus being more consistent with human perception.
Furthermore, to bridge the gap between the generative model and human perception, we introduce the perceptual loss [19], which captures the structurally correlated high-level features and thus is robust to noise and local artifacts. Specifically, we compute the perceptual loss $\mathcal{L}_{\text{p}}$ based on the $\mathrm{L}1$ distance of different convolutional layers within a pre-trained VGG-19 [37]:

### v3

@@ 1 Introduction
Visual realism with global illumination is a key demand of immersive virtual reality (VR) applications to intrigue users’ immersion arousal and respond realistically to virtual environments and events [yu2012visual]. However, realistic rendering of such globally-illuminated scenarios in real time remains a big challenge, especially in handling complex lighting situations like dynamic changes of illumination and environment lighting [mortensen2008real]. While photo-realistic rendering techniques (e.g., Monte Carlo path tracing) can synthesize images with global illumination, they are often extremely time-consuming and may take hours to render. Meanwhile, various works attempt to use approximations along with classical rasterization-based rendering, thus approximating global illumination effects at interactive or real-time performance. They generally operate in either image or 3D space and adopt different precomputation techniques to reduce runtime cost [sloan2002prt, keller1997instant]. Recently, learning-based techniques for photo-realistic rendering were proposed, some of which combine data-driven methods with classic rendering pipeline, resulting in improved image quality or rendering speedup [dai2020pbr, vicini2019learned]. These techniques, namely neural rendering [tewari2020state], usually provide stronger supervision to generative models by applying the prior knowledge in classical rendering techniques to neural networks.
Predicting indirect illumination is, nonetheless, a challenging task. Prior works generally regard this as a special case of image-to-image translation, and use convolutional neural networks (CNN) for indirect/global illumination prediction. The auxiliary geometry features are provided as inputs to predict global illumination, further conditioned on input the direct lighting [nalbach2017deep, xin2021lightweight] or emission maps [dai2020pbr]. However, several major challenges remain:
@@
To address these issues, we adopt a learning-based approach to address the issue of global illumination, following the principles of the rendering equation [kaj86re]. Our approach includes a learnable module that aggregates global features while utilizing geometry supervision. This module incorporates a modified multi-head attention mechanism [vaswani2017attention] to calculate the distribution based on geometry features, which guides the global aggregation of spatial features. This non-local process can be analogized to an integral over screen space, where the geometry-aware attention works as a learnable counterpart of the geometry term. Our proposed module overcomes the network bottleneck of the deep stack of convolutional layers, requiring less computation and fewer parameters while better capturing the long-range dependencies. Moreover, it can be applied to other tasks where geometry supervision is available and long-range dependency modeling is preferred.
Second, we present a monochromatic model design, where each RGB channel is independently processed using a shading generator conditioned on shared features from a geometry encoder. The redundant inter-channel interactions (introduced by the interconnections of convolutional filters) can be eliminated, and this relieves the burden of training data and network capacity, and enforces the network to learn the accurate color of indirect light. Furthermore, this makes our model immune to color variations, therefore allowing better generalization to novel and complex scenes and dynamic lighting, which are not present in the training data. Since HDR images can provide a much wider range of brightness and contrast for the high sensitivity of the human visual system to enhance the user’s perception of immersion and presence in the immersive virtual environment [GIHDR2008, HDRVR2022], our model is trained to synthesize HDR global illumination.
Finally, we created a novel synthetic HDR dataset of indoor scenes for network training based on the public indoor layout dataset 3D-Front [fu20203dfront]. Our dataset comes with well-annotated lighting, texture, and geometry features; the surfaces are also assigned with physically based materials, providing sufficient diversity for geometry features and visual appearances. We design an adversarial training framework to directly optimize our network on HDR data, where the discriminator and the perceptual loss are jointly used for network optimization.
@@ 2.1 Real-Time Global Illumination
Numerous works seek to acquire global illumination effects at a low time budget. Classical methods adopt different approximations to reduce performance costs. Some of them operate in 2D screen space, such as screen space ambient occlusion (SSAO) [bavoil2008image] and screen space directional occlusion (SSDO) [ritschel2009ssdo], using rasterized direct lighting and g-buffer features to approximate inter-reflections between surfaces. While some others operate in 3D space, usually achieving more realism with higher computational costs and storage, such as voxel cone tracing (VXGI) [crassin2011vxgi] and light propagation volumes (LPV) [kaplanyan2009lpv]. Lastly, precomputation-based methods like precomputed radiance transfer (PRT) [sloan2002prt] reduce the runtime cost by computing pre-integrations on the environment and BRDFs. Recently, with the advances in deep learning, data-driven methods have also been adopted to improve rendering quality efficiently. For example, training CNNs to denoise low sample-count images [chaitanya2017mc], or predicting the proper sample count for real-time raytracing [kuznetsov2018deep].
@@ 2.2 Image-to-Image Translation
The task of image-to-image translation converts images from one domain to another, e.g., image style translation, colorization, etc. The architecture of convolutional neural networks (CNN) has been the most commonly used technique for image-to-image tasks, such as U-Net [ronneberger2015u] and ResNet [he2016deep]. After the generative adversarial network (GAN) proposed by Goodfellow et al. [goodfellow2014generative], it has been further extended to conditional generative models and introduced to image-to-image tasks as well [isola2017image]. Recently, there has been another trend of using pre-trained models (e.g., VGG-19 [simonyan2014vgg]) to capture structurally correlated high-level image features for network optimization, namely the perceptual loss [johnson2016perceptual]. We show that a model’s output quality and robustness could be enhanced by jointly introducing adversarial loss and perceptual loss into network optimization.
@@ 2.3 Network Conditioning
Image-to-image translation could also be regarded as conditional synthesis, where networks learn to generate images conditioned on input geometry or semantic labels. Most of the early works simply provide conditioning data as the input of the network [isola2017image]. In contrast, many recent works focus on finding better ways to condition the image generation process, e.g., using a depth map to explicitly guide the convolution [wang2018depth], using learned transformations to modulate convolutional features [park2019semantic]. While these works have achieved great success, they are mainly used along with convolution layers, and thus remain local conditioning methods.
Inspired by the great success of attention mechanisms in modeling global dependencies [vaswani2017attention, zhang2019self] and motivated by our task formulation, we propose a feature aggregation method that models long-range dependencies guided by geometry features, where the aggregation weights are obtained from a modified multi-head attention mechanism. Different from the prior work on integrating attention to CNN [zhang2019self], we 1) explicitly use geometry features to guide the aggregation of global features; then 2) extend it into multi-head attention [vaswani2017attention], and find it sufficient to equip only bottleneck layers with attention. Both bring a significant reduction of computational overhead, with no observable performance degradation in our experiments.
@@ 2.4 Global illumination prediction
Several recent works use neural networks to predict global illumination given auxiliary shading and geometry features. Nalbach et al. [nalbach2017deep] first introduced CNN into global illumination prediction. They use a U-shaped CNN, taking pixel-wise features as inputs and predicting globally illuminated images. Bi et al. [bi2019deep] train GANs to predict global illumination, and Dai et al. [dai2020pbr] use the perceptual loss for network optimization. Then Xin et al. [xin2021lightweight] used a lightweight CNN and predicted the indirect component similar to the task setup in our work. However, Nalbach et al. [nalbach2017deep] used global illumination computed in screen space as their ground truth, and Dai et al. excluded textures in their ground truth data generation. Most prior works used LDR data for training or took no special care in training in linear HDR space. Moreover, they generally use stacked convolutional operations for shading learning, limiting the receptive fields to local neighborhoods, and thus it is difficult to predict global effects like indirect illumination.
@@ 3.1 Task Formulation
We aim to predict indirect lighting given direct lighting and auxiliary geometry features, including surface normal, depth, reflectance, etc. All these inputs can be efficiently computed using a rasterization-based rendering pipeline in real time. For simplicity, we assume all Lambertian surfaces within the scene, i.e., the BRDF term $f(\mathrm{p},\omega_{i},\omega_{o})$ is a constant for arbitrary $\omega_{i}$ and $\omega_{o}$. We first briefly review the concepts of rendering equation [kaj86re] and physically based rendering; by splitting the incident radiance $L_{i}$ into direct and indirect terms, we derive an alternative form of the surface rendering equation:
@@
where ${L}_{\mathrm{o}}^{\mathrm{d}}$ and ${L}_{\mathrm{o}}^{\mathrm{ind}}$ are the direct and indirect components of the scattering radiance, respectively. $G(\mathrm{p},\,\mathrm{p^{\prime}})=\frac{V(\mathrm{p^{\prime}}\leftrightarrow\mathrm{p}){\cos\theta\cos\theta^{\prime}}}{\parallel\mathrm{p}-\mathrm{p^{\prime}}\parallel}$ is the geometry term where $V(\mathrm{p^{\prime}}\leftrightarrow\mathrm{p})$ indicates the visibility between two vertices. We can obtain the direct lighting term ${L}_{\mathrm{o}}^{\mathrm{d}}$ by integrating over all light sources or by real-time alternatives [sloan2002prt, fernando2005pcss, heitz2016real]. We then feed ${L}_{\mathrm{o}}^{\mathrm{d}}$ into the network and predict the indirect component ${L}_{\mathrm{o}}^{\mathrm{ind}}$.
To apply fine-grained supervision to the network, we first decompose the direct and indirect components of the final illumination, letting the network focus on predicting the indirect component instead of global illumination. We then use reflectance factorization to demodulate shading and diffuse reflectance, thus effectively tasking the network with learning the demodulated shading (instead of scattered radiance), which has a relatively low frequency after multiple bounces within the scene. The final rendering is obtained by multiplying the predicted shading with the reflectance. Moreover, we propose a feature aggregation module for better modeling global dependencies, which is intuitively similar to an integral over screen space. We employ the modified attention mechanism [vaswani2017attention] to impose the effects of distance and visibility on network models, which intuitively works like a learnable counterpart of the geometry term, as is discussed later.
@@ 3.2 Model Design
Overview Our model contains a shading generation branch and a geometry encoding branch, both using a U-shaped network with skip-connections [ronneberger2015u]. We integrate 1) our feature aggregation module into the shading generator to model long-range dependencies; 2) the monochromatic design for a more consistent deep rendering process. We also utilize conditional modulation techniques [park2019semantic] for better geometry-aware supervision. The overview of our network architecture is shown in Fig. 2.
Geometry Encoder We use an additional geometry encoder that encodes geometry properties (i.e., positions $\mathbf{P}$, normals $\mathbf{N}$, and depth $\mathbf{D}$), leaving the other inputs to the shading encoder. The encoded geometry features are provided to condition the synthesis process of the shading generator. The separation of the encoding process splits the task into spatial structure modeling and light transport simulation, thus alleviating the requirements on training data and network complexity. Instead of simply concatenating the geometry features to the inputs of the generator layers, we use a more effective method, namely geometry-aware conditional modulation (GCM) similar to [park2019semantic]. Specifically, the conditional geometry features are transformed into scaling and biasing factors through learnable convolution layers as:
@@
Geometry-Aware Feature Aggregation Fully convolutional networks (FCN) will result in a limited-sized receptive field [xin2021lightweight, nalbach2017deep, dai2020pbr], i.e., the output pixel value is conditioned only on a local region. While deep stacks of convolutional operations can enlarge the receptive field, they also bring prohibitive computational costs, making network optimization difficult. In this work, we expect the network to generate indirect illumination from features of all locations within the image space. A learnable module is needed that can efficiently aggregate global features guided by geometry.
We use a modified multi-head attention mechanism [vaswani2017attention] to acquire the weights for the feature aggregation module based on geometry features (i.e., positions, depth, and surface normals). For each spatial location, the module first calculates the attention weights based on the encoded geometry features $f_{g}$, then obtains the response of that position as a weighted combination of the features from all locations. The attention weights $\mathrm{Attn}(j\,|\,i)$ for location $i$ are computed as:
@@
In addition, we find it beneficial to use multiple learnable linear transformations within attention computation, i.e., obtain multiple attention distributions $\mathrm{Attn}_{k}$ using different sets of linear transformations $Q_{k}$, $K_{k}$, and $V_{k}$, where k=1hk=1...h is the index of attention head and $h=8$ in our experiments. The aggregated features $\hat{x_{k}}$ are then concatenated to produce the final output $\hat{x}$, resulting in the multi-head version [vaswani2017attention] of the aggregation module. This allows the network to simultaneously attend to different subspaces of the spatial features, resulting in a better capacity for modeling long-range dependencies.
@@ 3.4 Adversarial Training and Loss Function
where $\omega_{\mathrm{c}}$, $\omega_{\mathrm{p}}$, and $\omega_{\mathrm{a}}$ are the hyperparameters controlling the weight of each loss term, we use empirical values where $\omega_{\mathrm{c}}=0.7$, $\omega_{\mathrm{p}}=0.28$, $\omega_{\mathrm{a}}=0.02$ in our experiments. We use smooth L1 loss [ren2015faster] as our content loss. While the L1 loss could not capture the inter-pixel correlations or the structural information of the image, we thus additionally introduced a PatchGAN [isola2017image] as the discriminator to obtain the adversarial loss term, which helps recover high-frequency local details. Moreover, it is an ill-posed problem to predict indirect illumination given direct lighting in screen space only (instead of in 3D solutions). While classic pixel-wise loss (e.g., L1, L2) generally tends to predict the average value over all possible solutions [ledig2017photo], we hope the adversarial training can guide the generator to approximate one particular solution, thus being more consistent with human perception.
Furthermore, to bridge the gap between the generative model and human perception, we introduce the perceptual loss [johnson2016perceptual], which captures the structurally correlated high-level features and thus is robust to noise and local artifacts. Specifically, we compute the perceptual loss $\mathcal{L}_{\text{p}}$ based on the $\mathrm{L}1$ distance of different convolutional layers within a pre-trained VGG-19 [simonyan2014vgg]: